A gas expands from 5 litre to 105 litre at a constant pressure 100N/m2. The work done is
A. 1 Joule B. 4 Joule C. 8 Joule D. 10 Joule

Grade:

1 Answers

Vikas TU
3 years ago
W=P*change in vol
chnge in vol=05-5=100litres=0.1m^3
W=100*0.1=10 joule
Was this answer helpful........................................................